OpenStudy (anonymous):

An inclined walkway leading to a new building is to rise 13 inches for each horizontal distance of 21 feet. Write this slope as a grade. (Round to the nearest tenth of a percent if necessary.)

OpenStudy (anonymous):

tan(\alpha) = 13/21

OpenStudy (anonymous):

so the grade \[\alpha\] is \[\alpha = \tan^{-1} (\alpha)\]

OpenStudy (anonymous):

a. 61.9% b. 1938.5% c. 5.2% d. 161.5% Those are the Choices =)

OpenStudy (anonymous):

my calculation gives 31.76°. so I am wrong... sorry....

OpenStudy (anonymous):

That's alright honey! Thanks =)